OpenStudy (anonymous):

solve x (16+2x)(10+8x)

OpenStudy (zzr0ck3r):

there is nothing to solve, when we solve for x, we are solving an equation. So I think you mean\[(16+2x)(10+8x)=0\] then we solve for x

OpenStudy (anonymous):

thank you

OpenStudy (zzr0ck3r):

in which case we have \(16+2x=0\implies2x=16\implies x=8\) does this make sense?

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):

ok now do the same thing for \(10+8x=0\)

OpenStudy (zzr0ck3r):

what do you get for \(x\) ?